# MSUHEP-24-019

### v1

@@ III Pseudo global analysis
where $M$, $Q_{T}$ and $Y$ are the invariant mass, transverse momentum and rapidity of the bosons, respectively. In this work, $M$ is set to 90 GeV, which is near the pole of the $W$ and $Z$ boson masses. Compared to $M$, the transverse momentum of the single vector boson production is usually small. Therefore $Q_{T}$ is ignored in this work. $\sqrt{s}$ is set to 13 TeV. To gain information for different $x$ values, the cross section observables are calculated with different $|Y|$ values of 0.1, 0.3, 0.5, 0.7, 0.9, 1.1, 1.3, 1.5, 1.7, 1.9, 2.1, 2.4, 2.8, 3.2, 4.0. The corresponding $x$ values range from 0.0001 to 0.38. In summary, 149 pseudo data points are introduced, corresponding to the DIS and Drell-Yan processes that are actually used in the PDF global analysis. For each data point, we assign a relative uncertainty of $3\%$. This is a precision comparable to the HERA measurements and the LHC Drell-Yan measurements. For experimental results other than HERA and LHC, such as earlier DIS and semi-inclusive DIS measurements, their precisions are much lower. The experiments actually used in the PDF global analysis are summarized in Ref. [2]. Although these pseudo observables are only rough estimations rather than real calculations, they reflect the complexcity of using multiple data inputs to determine the PDF shape parameters.
@@
With the constructed PDFs, the pseudo observables and the assigned uncertainty, the Hessian matrix can be calculated accordingly. Since there are 24 free parameters in this test, the same number of eigenvectors and corresponding eigenvalues are acquired by diagonalizing the Hessian matrix. The eigenvalues in this test are shown in Fig. 2. As explained in Ref. [1], the eigenvalues in logarithmic scale rank as a line, reflecting the sensitivities of the input data. An eigenvector with large eigenvalue usually indicates that the information related to this eigenvector is well constrained by the input data. Otherwise, the input data has poor sensitivity to this eigenvector, leading to a small eigenvalue. Although data points in this analysis are much more than the free parameters being determined and the assigned uncertainty of $3\%$ is quite small, there is still some PDF information for certain $x$ regions and parton flavors, for which no pseudo observable gives good sensitivity. Therefore, some eigenvalues are small compared to others. This also happens in a real PDF global analysis [2, 3, 7].
@@
A simple test can numerically show whether the data inputs in a global fit provide a strong constraint to a particular eigenvector, as to be illustrated below. Once the Hessian matrix is diagonalized and the orthogonal displacements $\{z_{i}\}$ are acquired, one can look into the $\chi^{2}$-$z_{i}$ relationship. Usually, if the eigenvector represented by $z_{i}$ is well constrained, $\chi^{2}$ is expected to be an ideal quadratic function of $z_{i}$, giving $\chi^{2}(z_{i}=\pm n;z_{j}=0$ for $j\neq i)=n^{2}$. In Fig. 3, the $\chi^{2}$-$z_{i}$ relationship for $i=22$ and $i=6$ are shown and compared, as an example. For $i=22$, the eigenvector has an eigenvalue at $\mathcal{O}(10^{4})$, thus its $\chi^{2}$-$z_{i}$ relationship follows a good quadratic approximation. For $i=6$ however, the eigenvalue is at $\mathcal{O}(1)$, resulting in a non-symmetric $\chi^{2}$-$z_{i}$ relationship. When the eigenvalue is small, the normalization factor $s_{k}\approx\sqrt{2/\epsilon_{k}}$ in Eq. (II) is enlarged. Therefore, any bias, such as inaccuracy in numerical calculations, would be significantly amplified. As a result, $z_{i}=\pm 1$ does not always yield $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$, especially for those poorly constrained eigenvectors. In conclusion, when the $\chi^{2}$-$z_{i}$ relationship does not follow a good quadratic approximation, the 1 standard deviation cannot be represented by $z_{i}=\pm 1$. Therefore, in this work, the one standard deviation in the uncertainty estimation of an observable is calculated with the $z_{i}$ value which yields $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$.
@@
Following the method described in the previous section, the first-order and second-order PDF error sets are acquired according to Eq. (9) and Eq. (11), respectively. The uncertainty of any physical observable $\mathcal{O}$ can be estimated accordingly. To numerically study the difference between the original Hessian method and the improved Hessian method, we compare the relative uncertainties on $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$, which are shown in Fig. 4. The improved Hessian method uncertainties, in which the second-order terms are included, are in general larger than the first-order calculations. Furthermore, the difference between the two uncertainties generally increases in the small and large $x$ region where the PDF uncertainty is larger. However, due to the complexity of the mixture of the parton information in multiple experimental observables, the uncertainties contributed from the second-order terms can still be relatively sizable, as compared to the first-order terms, even when the PDF uncertainty is small, i.e. when the PDFs are better constrained.
Figure 4: Relative uncertainties on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$, calculated by including only the first-order derivatives (dashed magenta curves) and with the addition of the second-order derivatives (solid red curves), respectively. The lower panel in each plot gives the ratio of the two uncertainties.
IV Pseudo global analysis with new data
The eigenvectors in a global analysis are determined from a particular set of data inputs. Therefore, both the first-order and second-order uncertainty estimations depend on the eigenvectors. To further test the improved Hessian method when the eigenvectors change, we introduce new experimental observables into the global fit, which are the new proton structure parameters $P_{u}$ and $P_{d}$ factorized from the forward-backward asymmetry of the Drell-Yan process [8], and the boost asymmetry $A_{\text{boost}}$ in the diboson process [9]. These asymmetry variables were recently proposed at hadron colliders which can provide unique information on the proton structure. The $P_{u}$ and $P_{d}$ parameters factorize the proton structure information in $pp(u\bar{u})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ and $pp(d\bar{d})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ events. The two light quark initial states can be decoupled because the electroweak couplings of $u$-$\bar{u}$-$Z$ and $d$-$\bar{d}$-$Z$ are different. As a result, $P_{u}$ and $P_{d}$ separately provide the $u$ and $d$ quark information, which are mixed and indistinguishable in the measurement of the total cross section $\sigma^{pp}_{Z}$. $A_{\text{boost}}$ is the asymmetry defined in the diboson production [9]. For example, in the $pp(u\bar{d})\rightarrow W^{+}\gamma\rightarrow\ell^{+}\nu+\gamma$ process, the $W^{+}$ boson predominantly couples to the $\bar{d}$ quark while the photon couples to the $u$ quark. Consequently, the boost asymmetry $A_{\rm boost}^{W^{+}\gamma}$ has a positive large value, namely $\gamma$ is more boosted than the $W^{+}$ boson, because the initial state parton lunminosity $u(x_{1}){\bar{d}}(x_{2})$ is larger than $u(x_{2}){\bar{d}}(x_{1})$ for $x_{1}>x_{2}$, etc. Therefore, the relative asymmetry between the initial states $q_{i}(u_{1})\bar{d}(x_{2})$ and $u(x_{2})\bar{d}(x_{1})$ ($x_{1}>x_{2}$), which arises from the nature of the valence and sea quark differences, can be measured by comparing the energy or rapidity of the two bosons.
$P_{u}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{u}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{u}(x_{1})$
$P_{d}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{d}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{d}(x_{1})$
$A^{W^{+}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{d}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{d}(x_{1})$
$A^{W^{-}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{u}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{u}(x_{1})$
To gain information on the $x$-dependence of the PDFs, the above pseudo observables are calculated with the $|Y|$ values the same as the previous settings. The relative uncertainty of each data point is also assigned as $3\%$. We repeat the calculations of the Hessian matrix now, using both the cross section observables and the new asymmetry observables; so we acquire the eigenvectors, eigenvalues, and the new orthogonal parameters $\{z_{i}\}$ accordingly. The new eigenvalues are shown in Fig. 5. With more data constraints introduced, the eigenvalues are in general larger. In Fig. 6, $\Delta\chi^{2}=\chi^{2}(z_{i}=1)-\chi^{2}_{\text{min}}$ is shown both in the previous pseudo global fit and the new fit. As discussed, $\Delta\chi^{2}=1$ indicates a good quadratic approximation, so that the corresponding eigenvector is well constrained by data. Therefore, more eigenvectors have $\Delta\chi^{2}=1$ or close to 1 in the new global fit, as expected.
Figure 5: 24 eigenvalues in the pseudo global analysis with new asymmetry observables included.
Figure 6: $\Delta\chi^{2}=\chi^{2}(z_{i}=1)-\chi^{2}_{\text{min}}$ in the pseudo global fit with new asymmetry observables included, compared with the previous global fit with only cross section observables.
The relative uncertainties on $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$ are shown in Fig. 7. As the constraints from the input data are strengthened, the estimated uncertainties are reduced with respect to the previous global fit in Fig. 4. In principle, when the uncertainties of individual eigenvectors decrease, the second-order uncertainties in Eq. (1) become less important with respect to the first-order ones. In Appendix B, we give another pseudo global fit in which the relative uncertainty of each data point is reduced to $0.1\%$. With such a high precision, there is indeed no difference between the first-order and second-order calculation, as expected. In the present section, although the new asymmetry observables provide stronger constraints, their precisions are not high enough to reduce the second-order terms to a negligible level. The ratios between the first-order and second-order uncertainties in Fig. 7 can become even larger than the previous global fit in Fig. 4, due to the complexity of the multiple data fit. This reflects the fact that the uncertainty estimation in a PDF global fit depends highly on the data inputs. Therefore, the second-order calculation is not only a necessary procedure to get a better uncertainty estimation, but also a sufficient way to evaluate the power of the data constraints in a global fit. When the difference between the first-order and second-order uncertainties is small, it indicates a strong data constraint so that the output PDF is less dependent on various assumptions and the choice of non-perturbative functions.
Figure 7: Relative uncertainty on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$. The new asymmetry observables $P_{u}$, $P_{d}$, and $A^{W^{\pm}\gamma}_{\text{boost}}$ are included in the global analysis.
V Conclusion
In this work, we propose an improved Hessian method in the PDF global analysis, which estimates the uncertainty with the inclusion of the complete set of second-order derivatives. A numerical test has been designed, in which a set of PDFs with multiple parameters are fitted to pseudo experimental observables. The test shows that when the precision of the input data used in a global analysis is limited, the improved Hessian method would give different uncertainties compared to the original method. Since the difference in the uncertainty can be reduced as the data constraints become stronger, one can tell whether a PDF global analysis is experimentally well constrained and less model-dependent, by comparing the uncertainties of the two methods.
@@ Acknowledgements.
Appendix A Appendix A: PDF-induced uncertainty on physical observables
In this section, we give detailed calculation of Eq. (14). Rewrite Eq. (1) as:
@@ Appendix A Appendix A: PDF-induced uncertainty on physical observables
Appendix B Appendix B: A pseudo global fit with high precision data
A pseudo global fit with high precision data is performed in this section. 9 data variables are used, including $\sigma^{\text{DIS}}_{\gamma}$, $\sigma^{\text{DIS}}_{Z}$, $\sigma^{pp}_{Z}$, $\sigma^{pp}_{W^{+}}$, $\sigma^{pp}_{W^{-}}$, $P_{u}$, $P_{d}$, $A^{W^{+}\gamma}_{\text{boost}}$ and $A^{W^{-}\gamma}_{\text{boost}}$. For each data point corresponding to a particular $x$ value, the relative uncertainty is set to $0.1\%$. With sych high precision data used in the global fit, the uncertainty of the second-order derivative calculation is consistent with the first-order calculation. The result is shown in Fig. 8.
Figure 8: Relative uncertainty on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$. The new asymmetry observables $P_{u}$, $P_{d}$, and $A^{W^{\pm}\gamma}_{\text{boost}}$ are included in the global analysis. Relative uncertainty of each data point is improved to $0.1\%$.

### v2

@@ Abstract
where $M$, $Q_{T}$ and $Y$ are the invariant mass, transverse momentum and rapidity of the bosons, respectively. In this work, $M$ is set to 90 GeV, which is near the pole of the $W$ and $Z$ boson masses. Compared to $M$, the transverse momentum of the single vector boson production is usually small. Therefore $Q_{T}$ is ignored in this work. $\sqrt{s}$ is set to 13 TeV. To gain information for different $x$ values, the cross section observables are calculated with different $|Y|$ values of 0.1, 0.3, 0.5, 0.7, 0.9, 1.1, 1.3, 1.5, 1.7, 1.9, 2.1, 2.4, 2.8, 3.2, 4.0. The corresponding $x$ values range from 0.0001 to 0.38.
We also introduce a group of new experimental observables into the global fit, which are the proton structure parameters, $P_{u}$ and $P_{d}$, factorized from the forward-backward asymmetry of the Drell-Yan process [11], and the boost asymmetry $A_{\text{boost}}$ in the diboson process [12]. These asymmetry variables were recently proposed at hadron colliders which can provide unique information on the proton structure. The $P_{u}$ and $P_{d}$ parameters factorize the proton structure information in $pp(u\bar{u})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ and $pp(d\bar{d})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ events. The two light quark initial states can be decoupled because the electroweak couplings of $u¯\bar{u}-$$Z$ and $d¯\bar{d}-$$Z$ are different. As a result, $P_{u}$ and $P_{d}$ separately provide the $u$ and $d$ quark information, which are mixed and indistinguishable in the measurement of the total cross section $\sigma^{pp}_{Z}$. $A_{\text{boost}}$ is the asymmetry defined in the diboson production [12]. For example, in the $pp(u\bar{d})\rightarrow W^{+}\gamma\rightarrow\ell^{+}\nu+\gamma$ process, the $W^{+}$ boson predominantly couples to the $\bar{d}$ quark while the photon couples to the $u$ quark. Consequently, the boost asymmetry $A_{\rm boost}^{W^{+}\gamma}$ has a positive large value, namely $\gamma$ is more boosted than the $W^{+}$ boson, because the initial state parton lunminosity $u(x_{1}){\bar{d}}(x_{2})$ is larger than $u(x_{2}){\bar{d}}(x_{1})$ for $x_{1}>x_{2}$, etc. Therefore, the relative asymmetry between the initial-state parton luminosities $u(x_{1})\bar{d}(x_{2})$ and $u(x_{2})\bar{d}(x_{1})$ ($x_{1}>x_{2}$) can be measured by comparing the energy or rapidity of the two bosons.
$P_{u}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{u}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{u}(x_{1}),$
$P_{d}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{d}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{d}(x_{1}),$
$A^{W^{+}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{d}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{d}(x_{1}),$
$A^{W^{-}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{u}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{u}(x_{1}).$
In summary, 149 pseudo data points are introduced, corresponding to the DIS and Drell-Yan processes that are actually used in the PDF global analysis. For each data point, we assign a relative uncertainty of $3\%$. This precision is comparable to the HERA measurements and the LHC Drell-Yan measurements. For experimental results other than HERA and LHC, such as earlier DIS and semi-inclusive DIS measurements, their precisions are much lower. The experiments actually used in the PDF global analysis are summarized in Ref. [2]. Although these pseudo observables are only rough estimations rather than real calculations, they reflect the complexity of using multiple data inputs to determine the PDF shape parameters.
@@
With the constructed PDFs, the pseudo observables and the assigned uncertainty, the Hessian matrix can be calculated accordingly. Since there are 24 free parameters in this test, the same number of eigenvectors and corresponding eigenvalues are acquired by diagonalizing the Hessian matrix. The eigenvalues in this test are shown in Fig. 3. As explained in Ref. [1], the eigenvalues in logarithmic scale rank as a line, reflecting the sensitivities of the input data. An eigenvector with large eigenvalue usually indicates that the information related to this eigenvector is well constrained by the input data. Otherwise, the input data has poor sensitivity to this eigenvector, leading to a small eigenvalue. Although data points in this analysis are much more than the free parameters being determined and the assigned uncertainty of $3\%$ is quite small, there is still PDF information for certain $x$ regions and parton flavors, for which no pseudo observable gives direct sensitivity. Therefore, some eigenvalues are small compared to others. This also happens in a real PDF global analysis [2, 3, 9].
@@
Once the Hessian matrix is diagonalized and the orthogonal displacements $\{z_{i}\}$ are acquired, one can look into the $\chi^{2}$-$z_{i}$ relationship. It is a good numerical test of the linear assumptions in the global analysis. Usually, if the eigenvector represented by $z_{i}$ is well constrained, $\chi^{2}$ is expected to be an ideal quadratic function of $z_{i}$, giving $\chi^{2}(z_{i}=\pm n;z_{j}=0$ for $j\neq i)=n^{2}$. As discussed before, it also means the data inputs related to this eigenvector $z_{i}$ are precise enough so that the corresponding $T_{\alpha}$ are approximately linear functions of the original parameters $\{a_{i}\}$ within small intervals. In Fig. 4, the $\chi^{2}$-$z_{i}$ relationship for $i=22$ and $i=6$ are shown and compared, as an example. $i=22$ is a typical eigenvector which has good linear approximation. Its eigenvalue is large at $\mathcal{O}(10^{4})$. The $\chi^{2}$-$z_{i}$ relationship follows a good quadratic approximation, so that $z_{i}=1$ yields $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$ corresponding to 1 standard deviation. For $i=6$ however, it is an eigenvector which is not well constrained by data, resulting in a non-symmetric $\chi^{2}$-$z_{i}$ relationship. In practice (also in this work), the 1 standard deviation in the uncertainty estimation is usually defined with the $\{z_{i}\}$ values which yields $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$.
@@
Following the method described in section IV, the uncertainty of any physical observable $\mathcal{O}$ can be estimated according to Eq. (19). To numerically study the difference between the original Hessian method and the improved Hessian method, we compare the relative uncertainties of $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$, which are shown in Fig. 5.
Figure 5: Relative uncertainties on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$. Comparison includes the contribution of the first-order derivatives which is estimated from the original Hessian method (blue curves), the uncertainty of $\delta\mathcal{O}_{E}$ which estimates the non-linearity of $T_{\alpha}(a_{i})$ (green curves), the uncertainty of the second-order derivatives additional to the first-order contributions (orange curves), and the total uncertainty from the improved Hessian method which is the quadratic sum of the previous three (red curves). The lower panel in each plot gives the ratio of the total uncertainty from the improved Hessian method to the original Hessian method.
Fig. 5 shows that the non-linear uncertainty in this pseudo global analysis is much larger than the original uncertainty, especially in the large and small $x$ region where no data can provide direct constraint. Even though the pseudo data inputs in this work are generated using the non-perturbative functions exactly the same as that used in the pseudo theory calculations, the uncertainty of the choice of non-perturbative functions can still be reflected in the non-constrained area. The DIS-liked observables are dominated by the valence quarks of $u_{V}$ and $d_{V}$ around the peak region of their distributions at $x\sim 0.1$, where the contributions of the sea quarks are relatively low. For the sea quarks however, there is no data predominated by them. Therefore, the valence pseudo quark PDFs are constrained more directly than the sea quark PDFs in this study. Consequently, the non-linear uncertainties for $u_{V}$ and $d_{V}$ are generally smaller than that for the sea quarks, as shown in Fig. 5. Moreover, the second-order uncertainty on the parton densities is not as significant as the non-linear uncertainty. If a physical observable contains information from multiple partons, the second-order uncertainties could be larger. In practice however, it could be fine not to compute the second-order uncertainty with thousands of second-order error sets for simplicity, especially when the non-linear uncertainty is already quoted.
In this work, the total uncertainties in the pseudo global analysis is several times larger than the original ones. According to the study from the CTEQ-TEA PDFs, the original uncertainties have to be enlarged corresponding to $\Delta\chi^{2}=37$, at the $68\%$ confidence level, to cover the difference of using various non-perturbative functions, which also leads to a total uncertainty several times larger. The MSHT20 PDF gives a similar conclusion. In their studies, the PDF uncertainty is defined by a dynamic tolerance which effectively corresponds to $\Delta\chi^{2}\sim 10$, giving the final uncertainty about 3 times larger than the original one. As a conclusion, the non-linear effect in the Hessian method is essential according to today’s experimental precisions.
@@ Acknowledgements.
Appendix A Appendix: PDF-induced uncertainty on physical observables
In this section, we give the detailed calculation of Eq. (17). Rewrite Eq. (6) as: